OpenStudy (haileyd):

By what factor does the average kinetic energy of gas molecules in a container increase when the temperature is raised from 300K to 900K

OpenStudy (rushwr):

\[E _{avg} = \frac{ 3 }{ 2 }nRT\] Where n = no. of moles R= universal gas constant T= temperature in Kelvins. Since they have not talked about a change in he no. of moles of gas n is constant and R is anyway constant. Therefore we can say \[E _{avg}\alpha T\] If Eavg = initial kinetic energy E'avg= final kinetic energy T' = final temperature T = initial temperature \[\frac{ E'_{avg} }{ E _{avg} } = \frac{ T' }{ T } = \frac{ 900 }{ 300 } = 3\]

OpenStudy (haileyd):

Thanks so much

OpenStudy (rushwr):

No problem bubs :)